Elmer's utility function is U(x, y) = min{x, y2}. If the price of x is $25 and the price of y is $15 and if Elmer chooses to consume 7 units of y, what must his income be?

1
Expert's answer
2021-11-17T10:03:18-0500

Given that

U(x, y) = min{x, y2}

Price of x is $25

ANd, the prcie of Y is $15

So,

25X + 15Y = M

if Y = 7,

So, 

At eqm, X = Y^​​​​​​2 = 49

Then ,

M = 25 × 49 + 15 × 7

= 1225 + 105

= 1330

Hence, the income is $1,330